# Cosmic-Ray Feedback from Supernovae in a Stratified Interstellar Medium

### v1

@@ 2.3 Energy Injections
We inject energy at random locations in the midplane. We restrict the midplane to be cells with vertical coordinate between $z=\pm H_{*}=\pm 100\,\mathrm{pc}$. To set the location of an injection, we select one of these cells from a random distribution proportional to the gas density in each cell, effectively following a Kennicutt-Schmidt law, $\mathrm{SFR}\propto\rho^{n}$, with $n=1$ (Kennicutt et al., 2007; Joung et al., 2009). While there is initially no positional preference in the $xy$ plane, the injections are more likely to occur inside dense gas at later times once density variations have developed.
@@
Once the number of injections and their locations are determined, we fill a cube with sides of length $L=100\,\mathrm{pc}$ centered on the injection location with a uniformly distributed thermal energy $E_{\mathrm{inj,th}}$ and uniformly distributed cosmic-ray energy $E_{\mathrm{inj,cr}}$. These energy values change between the CRInj and NoCRInj simulations (see Table 1). It is important to note this limits the accuracy of thermal feedback at the end of the simulation once a significant amount of cold gas has formed. The Sedov-Taylor phase ends at smaller radii for a supernova in a higher density medium, so the evolution of the thermal energy profile is inaccurate (see numerical implementations of energy injection like Kim & Ostriker (2015) for solutions to this problem). As a result, we underestimate the total thermal feedback, in particular the conversion to momentum feedback, towards the end of the simulation.
This caveat does not apply as significantly to the cosmic-ray feedback. Since the cosmic-ray energy is predominantly diffusive in the high density regime (because of decreased Alfven speed), the cosmic-ray energy quickly diffuses along the local magnetic field. However, our $10\,\mathrm{pc}$ resolution could limit the accuracy of the local magnetic field structure around energy injection sites. A more resolved magnetic field could amplify the cosmic-ray feedback on small scales. Additionally, we do not include cosmic-ray momentum boost of supernovae explosions (Diesing & Caprioli, 2018), or a local decrease in the cosmic-ray diffusion coefficient in star formation regions (Semenov et al., 2021). These additional complexities in cosmic-ray transport would primarily change cosmic-ray feedback in the midplane, but would have a smaller impact on larger length scales $(L\gtrsim 1\,\mathrm{kpc})$ and longer timescales $(t>1\,\mathrm{Myr})$ we consider.
@@ 2.4 A Note on Parker Instability
Our initial conditions are overall similar to setups used to examine the Parker instability (Parker, 1966). We tested multiple setups in the single injection simulations and found disruption by the injection is dependent on the Parker stability criterion (Habegger et al., 2023). Our initial setup in this work (see Table 1) is linearly unstable to the Parker instability. Since we begin with an isothermal atmosphere, the linear instability criterion (only considering a plane-parallel atmosphere) is (Newcomb, 1961; Zweibel, 2017; Heintz & Zweibel, 2018):
@@
and plugging in the values from Table 1, we find $\gamma_{c}<\frac{4}{3}$. Since we include non-advective transport like streaming and diffusion, we expect to satisfy this criteria and the setup is then unstable.
We can also estimate the growth rate of the most unstable mode. Following the dispersion relations derived in Heintz & Zweibel (2018), we use the 3D Modified Parker (non-zero $\gamma_{c}$) with streaming transport case. For our setup, we find the most unstable Fourier mode has a wavelength along the magnetic field of $L_{x}\sim 0.85\,\mathrm{kpc}$, and a growth rate of $\tau\sim 33\,\mathrm{Myr}$. Additionally, all the the wavelengths above $L_{x}\geq 0.35\,\mathrm{kpc}$ are unstable. Those large scale wavelengths are all above the size of our injections, which are a box of side length $L=100\,\mathrm{pc}$ (see Section 2.3).
@@ 3 Results
We now describe the results of our two simulations, named CRInj and NoCRInj. A key similarity between the simulations is the development of two stages in time. We describe this time evolution before a detailed examination of each stage. The first stage is related to the exponential growth of the Parker instability, while the second stage is a steady state which develops after the Parker instability has disrupted the initial hydrostatic equilibrium. Finally, we compare our simulations to observational diagnostics.
Figure 3: The time evolution of energy density averaged over the simulation volume in simulations CRInj (left column) and NoCRInj (right column). The top row shows the kinetic energy density divided into its directional components $(\hat{x},\hat{y},\hat{z})$, and the middle row shows the same for magnetic energy density. The bottom row compares thermal energy density, total magnetic energy density, total kinetic energy density, and cosmic-ray energy density.
Figure 4: Top: Mass outflow and inflow rates for regions above and below the midplane $(|z|>1\,\mathrm{kpc})$. Solid lines show the CRInj simulation and dashed lines show the NoCRInj simulation. Bottom: Mass fraction of different phases in terms of the initial mass in the simulation. After $t=0$, the sum of these phases can be less than one because mass can flow out of the top and bottom boundaries. Again, the solid lines show the CRInj simulation and the dashed lines show the NoCRInj simulation.
@@ 3.1 Time Evolution
We run the simulations for $400\,\mathrm{Myr}$. Initially, the entire setup is destabilized by many energy injections. These perturbations drive the Parker instability, producing a fountain flow which begins at approximately $100\,\mathrm{Myr}$ into the simulation. At $200\,\mathrm{Myr}$ the simulation begins to settle into a steady state structure. This steady state lasts from approximately $250\,\mathrm{Myr}$ until the end of the simulation at $400\,\mathrm{Myr}$.
@@
These evolutionary stages are apparent in Figure 3, which illustrates how the different components of energy density evolve over time. The left column of Figure 3 shows the results from the CRInj simulation, and the right column shows the results from the NoCRInj simulation.
@@
In the top row of Figure 3, we show the directional components of kinetic energy density for each simulation. During the initial destabilization, most of the motion ends up being vertical, perpendicular to the midplane. Before $t\sim 250\,\mathrm{Myr}$, the CRInj simulation produces more kinetic energy in every component direction than the NoCRInj simulation. Therefore, the additional cosmic-ray pressure from the injections has a significant impact on the dynamics of the gas. In the eventual steady state after $t\sim 250\,\mathrm{Myr}$ the kinetic energy in each component levels off to a constant value. The kinetic energy parallel to the mean magnetic field ($\hat{x}$ direction) ends up being the largest. The spikes are due to individual energy injections, which quickly dissipate. A key difference between the simulations in the steady state is the ordering. In the CRInj simulation, the vertical kinetic energy is distinctly larger than the planar motion perpendicular to the magnetic field ($\hat{y}$ direction). However, the NoCRInj simulation has these energies at about the same value.
@@
The middle row of Figure 3 shows the time evolution of the directional components of magnetic energy density. Both simulations have a similar evolution - the magnetic field starts only in the $\hat{x}$ direction, and this energy decays because magnetic flux in the $\hat{x}$ and $\hat{y}$ direction can escape out of the top and bottom of the simulation box. The magnetic energy density in the vertical direction increases, exponentially, until approximately $t\sim 180\,\mathrm{Myr}$. In Section 3.2, we show this increase is tied to the Parker instability’s characteristic structure. After $t\sim 180\,\mathrm{Myr}$, the magnetic energy density decays. This decay might be offset by amplification due to shearing if we had included galactic rotation in our models. Additionally, we observe numerical magnetic reconnection, which contributes to the loss of magnetic energy density.
@@
The bottom row of Figure 3 compares the evolution of total thermal, magnetic, kinetic, and cosmic-ray energy density. The total kinetic energy density is much lower than the cosmic-ray, magnetic, and thermal energy densities. The thermal energy density has continual spikes as a result of the energy injections, but an underlying decrease as cold, dense gas forms from optically thin radiative cooling. The total magnetic energy reflects the decrease in the $\hat{x}$ directed component and the spike in vertically directed magnetic field. Finally, the cosmic-ray energy density shows the biggest difference between the two simulations. The CRInj simulation shows a linear increase early on as the cosmic-ray energy injections add to the background cosmic-ray energy density. Both simulations show a decrease in cosmic-ray energy density which is tied to the increase in vertical magnetic field. Since the cosmic rays are stuck to the local magnetic field, there needs to be magnetic flux directed out of the top and bottom of the simulation to allow the cosmic-ray energy to escape. As the vertical magnetic field energy density decreases, the cosmic-ray energy density levels off in the NoCRInj simulation. The CRInj simulation also levels off, but increases slightly due to continuing cosmic-ray energy injections.
@@
Figure 5: Left: Kinetic energy spectrum $\tilde{v}_{x}^{2}(k)$ (in units of $\mathrm{km}^{2}\,\mathrm{s}^{-2}$) for the CRInj simulation at each $1\,\mathrm{Myr}$ time dump. The color map shows the amplitude of velocity fluctuations at each wavelength and the green line traces the peak wavelength of the spectrum. Right: Amplitude over time of the four longest wavelength Fourier modes, along with a line showing the maximum growth rate predicted by linear theory.
@@
Moving on from energy evolution, we focus on mass evolution in the simulations. The top graph of Figure 4 shows the mass inflow and outflow rates outside of the midplane $(|z|>1\,\mathrm{kpc})$. The solid lines show the CRInj simulation and the dashed lines show the NoCRInj simulation. The cosmic-ray injections lead to a larger peak mass outflow (and inflow) rate, and drive a larger steady outflow. In the NoCRInj simulation, the inflow and outflow nearly balance each other at late times $(t>300\,\mathrm{Myr})$. There also appears to be a fountain flow component, where outflows are followed by inflows at a delay of about $\sim 30\,\mathrm{Myr}$ at the start of the simulations. Finally, the two simulation curves are almost exactly the same until $t\sim 100\,\mathrm{Myr}$, which reflects that the simulations were run with the same random seed.
@@
In the bottom of Figure 4, we show the mass fractions of different gas phases. We define warm gas as gas with temperature $4\cdot 10^{3}\mathrm{K}\lesssim T\lesssim 10^{5}\mathrm{K}$, unstable gas as gas with temperature $100\mathrm{K}\lesssim T\lesssim 4\cdot 10^{3}$, and cold gas is gas with temperature $T\lesssim 100\mathrm{K}$. In pressure-density phase space (see Figure 2) the warm gas stretches between the minimum of the equilibrium curve up to the peak above a gas density of $1\,\mathrm{cm}^{-3}$. The unstable gas is between that peak and the minimum just below a density of $10\,\mathrm{cm}^{-3}$, and the cold gas ranges from that minimum to the temperature floor of $T=30\mathrm{K}$. It should be noted that the lines in this plot do not necessarily add to unity at each time. There is a small loss of mass out of the outflow boundaries which causes the total mass to decrease by a factor of $\lesssim 3\%$ over the $400\,\mathrm{Myr}$ of evolution.
@@
Overall, we see that initially $(t\lesssim 120\,\mathrm{Myr})$ some thermally unstable gas is created, which eventually becomes cold gas. The lack of a significant mass fraction of unstable gas is characteristic of weak multiphase turbulence (Ho et al., 2024). That observation is consistent with our simulations - most of the kinetic energy in our simulations are vertical, and associated with the bulk outflow in the top panel of Figure 4.
@@
After the unstable gas forms, the cold gas quickly becomes the dominant component of the overall mass fraction after $t\sim 300\,\mathrm{Myr}$ for both simulations. There is a slight delay in the NoCRInj simulation, which is shown with dashed lines. The peak in unstable gas formation tracks with the inflow rate in the top plot of Figure 4. This correlation suggests a connection between the fountain flow and the formation of cold gas via thermal instability. In Section 3.2, we show this connection comes from a combination of the Parker instability and Field’s thermal instability.
@@
Figure 6: Left: Vertical velocity colormap with gas density contours and magnetic field lines overlayed. The large plume centered at $x\sim 0.1kpc$ is moving upward at a fast rate $(v_{z}\sim 50-100\,\mathrm{km}\,\mathrm{s}^{-1})$. Right: Colormap of cosmic-ray pressure with gas density contours and magnetic field lines overlayed. The density contours highlight the thermally unstable gas phase which forms below where the plume expands into itself (because of periodic boundary conditions). The plume is driven upward partly by the cosmic-ray pressure gradient, which is apparent in the colormap. Both the left and right simulation slices are taken from the CRInj simulation at $y=-395\mathrm{pc}$ and at a time of $t=165\,\mathrm{Myr}$.
Figure 7: Top: Growth of vertical kinetic and magnetic energy density in the CRInj simulation is exponential with a characteristic time of $\tau\sim 40\,\mathrm{Myr}$. Middle: Column density colormap and average cosmic-ray pressure contours at $t=100\,\mathrm{Myr}$. Bottom: Same as middle plot, but at $t=150\,\mathrm{Myr}$.

### v2

@@ 2.3 Energy Injections
We inject energy at random locations in the midplane. We restrict the midplane to be cells with vertical coordinate between $z=\pm H_{*}=\pm 100\,\mathrm{pc}$. We select a random $x,y,z$ position following that restriction. This random selection neglects the fact that star formation occurs in dense gas (Schmidt, 1959). However, it is a reasonable choice; others have found that injecting supernovae primarily in dense gas can lead to peak-driving which reduces the impact they have when they occur in low-density gas, and thus fails to create a realistic ISM (Walch et al., 2015).
@@
It is important to note that we are considering a low rate of supernovae, $\dot{N}_{\mathrm{SN}}=1\mathrm{Myr}^{-1}$ in comparison to other tallbox simulations. This rate approximately corresponds to a star formation rate of $100M_{\odot}\mathrm{Myr}^{-1}\mathrm{kpc}^{-2}=10^{-4}M_{\odot}\mathrm{yr}^{-1}\mathrm{kpc}^{-2}$. This rate is reasonable in the $\sim\mathrm{kpc}^{2}$ area near the solar neighborhood, between spiral arms (Elia et al., 2022; McCallum et al., 2025).
Once the number of injections and their locations are determined, we inject thermal energy equally in a sphere of radius $20\,\mathrm{pc}$. We also inject kinetic energy (with velocity directed radially outward from the injection location) following the prescription given by Equation 34 in Kim & Ostriker (2015). Finally, we inject cosmic ray energy in a sphere of radius $50\,\mathrm{pc}$ centered on the same location. Both simulations we present inject the same kinetic energy with each supernova, but we vary the thermal energy and cosmic ray energy injected. In CRBkg, we only inject thermal energy, and in CRInj we inject some cosmic ray energy (see Table 1).
@@ 2.4 A Note on Parker Instability
Our initial conditions are overall similar to setups used to examine the Parker instability (Parker, 1966). We tested multiple setups in the single injection simulations and found disruption by the injection only occurs for setups unstable to the Parker instability (Habegger et al., 2023). Our initial setup in this work (see Table 1) is linearly unstable to the Parker instability. Since we begin with an isothermal atmosphere, the linear instability criterion (only considering a plane-parallel atmosphere) is (Newcomb, 1961; Zweibel, 2017; Heintz & Zweibel, 2018):
@@
and plugging in the values from Table 1, the instability criterion becomes $\gamma_{c}<\frac{4}{3}$. Since we include non-advective transport like streaming and diffusion, we expect to satisfy this criteria and the setup is then unstable to the Parker instability.
We can also estimate the growth rate of the most unstable mode. Following the dispersion relations derived in Heintz & Zweibel (2018), we use the 3D Modified Parker (non-zero $\gamma_{c}$) with streaming transport case. For our setup, we find the most unstable Fourier mode has a wavelength along the magnetic field of $L_{x}\sim 0.81\,\mathrm{kpc}$, and a growth rate of $\tau\sim 32\,\mathrm{Myr}$. Additionally, all the the wavelengths above $L_{x}\geq 0.32\,\mathrm{kpc}$ are unstable. Those large scale wavelengths are all above the characteristic size of a fully evolved supernova supernova remnant, which generally expand to be $100\,\mathrm{pc}$.
@@ 3 Results
We now describe the results of our two simulations, named CRInj and CRBkg. A key similarity between the simulations is the development of two stages in time. We describe this time evolution before a detailed examination of each stage. The first stage is related to the exponential growth of the Parker instability, while the second stage is a steady state which develops after the Parker instability has disrupted the initial hydrostatic equilibrium. Finally, we compare our simulations to observational diagnostics.
Figure 3: The time evolution of simulations CRInj (left column) and CRBkg (right column), with each quantity averaged over the entire simulation volume. The top row shows the average velocity divided into its directional components $(\hat{x},\hat{y},\hat{z})$, and the middle row shows the same for magnetic field strength. Compared to the simulation with only an initial background of cosmic rays, additional cosmic ray injections amplify the vertical velocity. The exponential growth of vertical magnetic field is related to the Parker instability (see Section 3.2). The bottom row compares thermal, magnetic, turbulent, and cosmic-ray pressures. The average cosmic ray pressure reaches a steady state near the kinetic pressure and below the average gas and magnetic pressures.
Figure 4: Top: Mass outflow and inflow rates for regions above and below the midplane $(|z|>1\,\mathrm{kpc})$. Solid lines show the CRInj simulation and dashed lines show the CRBkg simulation. The outflows and inflows are out of phase, suggesting the simulations produce a fountain flow. Since the effective star formation rate in our simulation volume is $\sim 10^{-4}M_{\sun}\mathrm{yr}^{-1}\mathrm{kpc}^{-2}$, the outflows have mass loading factors $\eta=\dot{M}_{\mathrm{out}}/\dot{M}_{\mathrm{SFR}}>10^{3}$. Bottom: Mass fraction of different phases in terms of the initial mass in the simulation. After $t=0$, the sum of these phases can be less than one because mass can flow out of the top and bottom boundaries. Again, the solid lines show the CRInj simulation and the dashed lines show the CRBkg simulation.
@@ 3.1 Time Evolution
We run the simulations for $400\,\mathrm{Myr}$. Initially, the entire setup is destabilized by many energy injections. These perturbations drive the Parker instability, immediately producing a fountain flow which becomes nonlinear at approximately $100\,\mathrm{Myr}$ into the simulation. At $200\,\mathrm{Myr}$ the simulation begins to settle into a steady state structure. This steady state lasts from approximately $250\,\mathrm{Myr}$ until the end of the simulation at $400\,\mathrm{Myr}$.
@@
These evolutionary stages are apparent in Figure 3, which illustrates how the average velocity, magnetic field strength, and pressure components in the simulation vary over time. The left column of Figure 3 shows the results from the CRInj simulation, and the right column shows the results from the CRBkg simulation.
@@
In the top row of Figure 3, we show the directional components of velocity averaged over the simulation volume. During the initial destabilization, most of the motion ends up being vertical, perpendicular to the midplane. Before $t\sim 250\,\mathrm{Myr}$, the CRInj simulation produces more motion in every component direction than the CRBkg simulation, but most significantly in the vertical direction. Therefore, the additional cosmic-ray pressure from the injections has a significant impact on the dynamics of the gas. In the eventual steady state after $t\sim 250\,\mathrm{Myr}$ the average velocity in each direction levels off to a constant value. The velocity parallel to the mean magnetic field ($\hat{x}$ direction) ends up being the largest. The spikes are due to individual energy injections. A key difference between the simulations in the steady state is the ordering. In the CRInj simulation, the vertical motion is distinctly larger than the planar motion perpendicular to the magnetic field ($\hat{y}$ direction). However, the CRBkg simulation has these velocities at about the same value.
@@
The middle row of Figure 3 shows the time evolution of the directional components of magnetic field averaged over the entire simulation volume. Both simulations have a similar evolution - the magnetic field starts only in the $\hat{x}$ direction, and this energy decays because magnetic flux in the $\hat{x}$ and $\hat{y}$ direction can escape out of the top and bottom of the simulation box. The magnetic field strength in the vertical direction increases, exponentially, until approximately $t\sim 180\,\mathrm{Myr}$. In Section 3.2, we show this increase is tied to the Parker instability’s characteristic structure. After $t\sim 180\,\mathrm{Myr}$, the magnetic field decays. This decay might be offset by amplification due to shearing if we had included galactic rotation in our models. Additionally, we observe numerical magnetic reconnection, which contributes to the decrease in magnetic field strength. Overall, the loss in horizontal magnetic field illustrates how the Parker instability can efficiently decrease, redistribute, and reorient the magnetic field
@@
The bottom row of Figure 3 compares the evolution of thermal, magnetic, turbulent, and cosmic-ray pressures averaged over the entire simulation volume. Initially, the midplane gas, magnetic, and cosmic ray pressures are all $10^{4}k_{B}\mathrm{K}\mathrm{cm}^{-3}$ (see Table 1). The total turbulent pressure is initially zero, but it rises to the order of magnitude of the other pressure components. The thermal and turbulent pressures have continual spikes as a result of the injections. The thermal pressure also has an underlying decrease as cold, dense gas forms from optically thin radiative cooling. The total magnetic pressure reflects the decrease in the $\hat{x}$ directed component and the spike in vertically directed magnetic field. Finally, the cosmic-ray pressure shows the biggest difference between the two simulations. The CRInj simulation shows a linear increase early on as the cosmic-ray energy injections add to the background cosmic-ray pressure. Both simulations show a decrease in cosmic-ray pressure which is tied to the increase in vertical magnetic field. Since the cosmic rays move along the direction of the local magnetic field, there needs to be magnetic flux directed out of the top and bottom of the simulation to allow the cosmic-ray energy to escape. As the vertical magnetic field strength decreases, the cosmic-ray pressure levels off in the CRBkg simulation. The CRInj simulation also levels off, but to a non-zero value due to continuing cosmic-ray energy injections.
@@
Figure 5: Left: Kinetic energy spectrum $\tilde{v}(k)$ (in units of $\mathrm{km}\,\mathrm{s}^{-1}$) for the CRInj simulation at each $1\,\mathrm{Myr}$ time dump. The color map shows the amplitude of velocity fluctuations at each wavelength $L=2\pi/k$ and the green line traces the peak wavelength of the spectrum. The spectrum is primarily driven by supernovae, at a scale of $L\sim 20\mathrm{pc}$. Eventually, the larger scale $L\gtrsim 0.33\mathrm{kpc}$ mode dominates. Right: Amplitude over time of four key Fourier modes, along with line showing the growth rate of the $L\gtrsim 0.5\mathrm{kpc}$. The $L=0.5\mathrm{kpc}$ mode grows as a result of the Parker instability, independent of the large scale modes until it reaches their level. While that mode never becomes dominant, it feeds the smaller scale modes once it grows to the same strength at $t\sim 80\,\mathrm{Myr}$, causing the $L\gtrsim 0.33\mathrm{kpc}$ mode to dominate. While the $L\gtrsim 0.5\mathrm{kpc}$ mode grows consistently until joining the other modes, the small scale modes saturate after $t\sim 20\,\mathrm{Myr}$.
@@
Moving on from energy evolution, we focus on mass evolution in the simulations. The top graph of Figure 4 shows the mass inflow and outflow rates outside of the midplane $(|z|>1\,\mathrm{kpc})$. The solid lines show the CRInj simulation and the dashed lines show the CRBkg simulation. The cosmic-ray injections lead to a larger peak mass outflow (and inflow) rate, and they drive a larger steady outflow at late times $(t>300\,\mathrm{Myr})$. In the CRBkg simulation, the inflow and outflow nearly balance each other at late times. There also appears to be a fountain flow component, where outflows are followed by inflows at a delay of about $\sim 30\,\mathrm{Myr}$. Finally, the two simulation curves are almost exactly the same until $t\sim 100\,\mathrm{Myr}$, when the Parker instability begins to change the magnetic field orientation.
@@
In the bottom of Figure 4, we show the mass fractions of different gas phases. Warm gas has a temperature $4\cdot 10^{3}\mathrm{K}\lesssim T\lesssim 10^{5}\mathrm{K}$, whereas unstable gas has a temperature $100\mathrm{K}\lesssim T\lesssim 4\cdot 10^{3}$. Cold gas has a temperature $T\lesssim 100\mathrm{K}$. In pressure-density phase space (see Figure 2) the warm gas stretches between the minimum of the equilibrium curve up to the peak above a gas density of $1\,\mathrm{cm}^{-3}$. The unstable gas is between that peak and the minimum just below a density of $10\,\mathrm{cm}^{-3}$, and the cold gas ranges down to the temperature floor of $T=30\mathrm{K}$. It should be noted that the mass fractions in Figure 4 do not add to unity at all times. There is a small loss of mass out of the outflow boundaries which causes the total mass to decrease by a factor of $\lesssim 3\%$ over the $400\,\mathrm{Myr}$ of evolution.
@@
Overall, we see that initially $(t\lesssim 120\,\mathrm{Myr})$ some thermally unstable gas is created, which eventually becomes cold gas. The lack of a significant mass fraction of unstable gas is characteristic of weak multiphase turbulence (Ho et al., 2024). That conclusion is consistent with our simulations - most of the kinetic energy in our simulations are vertical, and associated with the bulk outflow in the top panel of Figure 4. Driving the gas into the strong turbulence regime would require more supernova injections.
@@
After the unstable gas forms, the cold gas quickly becomes the dominant component of the overall mass fraction after $t\sim 250\,\mathrm{Myr}$ for both simulations. There is a slight delay in the CRBkg simulation, which is shown with dashed lines. The peak in unstable gas formation tracks with the inflow rate in the top plot of Figure 4. This correlation suggests a connection between the fountain flow and the formation of cold gas via thermal instability. In Section 3.2, we show this connection comes from a combination of the Parker instability and Field’s thermal instability.
@@
Figure 6: Left: Vertical velocity colormap with gas density contours and magnetic field lines overlaid. The large plume centered at $x\sim 0.1kpc$ is moving upward at a fast rate $(v_{z}\sim 50-100\,\mathrm{km}\,\mathrm{s}^{-1})$. Right: Colormap of cosmic-ray pressure with gas density contours and magnetic field lines overlaid. The density contours highlight the thermally unstable gas phase which forms below where the plume expands into itself (because of periodic boundary conditions). The plume is driven upward partly by the cosmic-ray pressure gradient, which is apparent in the right side plot. Both the left and right simulation slices are taken from the CRInj simulation at $y=352.5\mathrm{pc}$ and at a time of $t=135\,\mathrm{Myr}$. The crimson box in the right hand figure shows the zoom-in region pictured in Figure 7.